### v1

# MalAlgoQA: A Pedagogical Approach for Evaluating Counterfactual Reasoning Abilities

Abstract
This paper introduces MalAlgoQA, a novel dataset designed to evaluate the counterfactual reasoning capabilities of Large Language Models (LLMs) through a pedagogical approach. The dataset comprises mathematics and reading comprehension questions, each accompanied by four answer choices and their corresponding rationales. We focus on the incorrect answer rationales, termed "malgorithms", which highlights flawed reasoning steps leading to incorrect answers and offers valuable insights into erroneous thought processes. We also propose the Malgorithm Identification task, where LLMs are assessed based on their ability to identify corresponding malgorithm given an incorrect answer choice. To evaluate the model performance, we introduce two metrics: Algorithm Identification Accuracy (AIA) for correct answer rationale identification, and Malgorithm Identification Accuracy (MIA) for incorrect answer rationale identification. The task is challenging since state-of-the-art LLMs exhibit significant drops in MIA as compared to AIA. Moreover, we find that the chain-of-thought prompting technique not only fails to consistently enhance MIA, but can also lead to underperformance compared to simple prompting. These findings hold significant implications for the development of more cognitively-inspired LLMs to improve their counterfactual reasoning abilities, particularly through a pedagogical perspective where understanding and rectifying student misconceptions are crucial. The data and code can be found at https://github.com/luffycodes/MalAlgoQA-Dataset.
1 Introduction
Counterfactual reasoning refers to the ability to consider alternative scenarios and infer the cause of an event based on hypothetical situations. It is a fundamental aspect of human cognition that plays a crucial role in problem-solving, decision-making, and learning [5, 13, 14]. Although large language models (LLMs) have demonstrated remarkable performance in various natural language processing tasks, their ability to engage in counterfactual reasoning remains relatively underexplored. Counterfactual reasoning requires a degree of hypothetical and abstract thinking, which poses a significant challenge for current LLMs, as they are generally trained to predict outcomes based on facts, rather than imagining alternative scenarios that might not be explicitly present in the training data [3, 9, 19].
Question Each side of a square is 3 inches long. A student increases the length of each side by 3 inches. The area of the new square is how many times the area of the original square?
Answer Choices (A) 9 (Rationale: Squared the change in the length.) (B) 2 (Rationale: Calculated the relationship between the sides.) (C) 3 (Rationale: Used the change in the length.) (D) 4 (Rationale: Calculated the area of the original square: $3\times 3=9$ and the area of the new square with each side length increased by 3 inches: $6\times 6=36$. Then determined that the area of the new square is 4 times the area of the original square.)
Table 1: An example question from the MalAlgoQA dataset illustrating the Malgorithm Identification task. Each answer choice is associated with a rationale, representing the reasoning process that led to that answer. The task for the model is to correctly identify the rationale given a particular answer choice. For incorrect answer choices, the corresponding rationales are malgorithms, representing flawed reasoning processes and thereby evaluating the counterfactual reasoning abilities of LLMs.
To fill this gap, we introduce MalAlgoQA, a novel dataset designed to evaluate the counterfactual reasoning capabilities of LLMs through a pedagogical approach with multiple-choice question answering. MalAlgoQA consists of a diverse set of 807 mathematics and 290 reading comprehension questions spanning various grade levels (3-11), content classifications (e.g., algebra, geometry, number & operations), and Depth of Knowledge levels (DOK) (1-3) [25]. Each question is accompanied by a set of answer choices and their associated rationales, which we term "malgorithms" (inspired by the concept of "mal-rules" proposed by [11]). Malgorithms represent the flawed reasoning that leads to an incorrect answer, revealing the underlying thought process that causes a student to arrive at the incorrect conclusion.
With MalAlgoQA, we propose the Malgorithm Identification task, which involves identifying the rationale behind the selection of a particular answer choice. When the given answer choice is correct, a model needs to employ causal reasoning to determine the rationale or algorithm that justified the correct selection. On the other hand, when the given answer choice is incorrect, the task demands counterfactual reasoning, as the model must identify the flawed reasoning or malgorithm that led to the selection of an incorrect option. In this case, the model must consider hypothetical scenarios and reason about the cause of an incorrect choice, which may not align with the facts. To illustrate the Malgorithm Identification task, consider a simple question: "What is 1 + 2 × 3 + 4?" with choices A. 11, B. 13, and C. 21. The rationales are: "Worked left to right" (corresponding to choice B), "Applied PEASMD rule instead of PEMDAS" (corresponding to choice C), and "Applied PEMDAS rule" (corresponding to the correct choice A). Given an answer choice, the model should identify the rationale explaining that choice.
In order to evaluate model performance, we introduce two key metrics: Algorithm Identification Accuracy (AIA), which measures accuracy in identifying the rationale behind a given correct answer. Malgorithm Identification Accuracy (MIA), of particular interest, evaluates the model’s counterfactual causal reasoning abilities by quantifying accuracy in identifying the rationale behind an incorrect answer. Our experiments reveal that GPT-4o [10], a state-of-the-art LLM, achieves an AIA of $95.7\%$, demonstrating its proficiency in this task. However, when it comes to MIA, GPT-4o’s performance drops significantly, with an accuracy of only $66.1\%$. This pattern, where accuracy is higher for correct answer rationales and lower for incorrect answer rationales, is consistent across other language models, such as LLaMA-3-70B [1] and GPT-3.5 [18]. It emphasizes the challenge of counterfactual reasoning for LLMs, especially on finding the underlying rationale behind incorrect answer choices.
Our experiments also reveal an unexpected result that Chain-of-Thought (CoT) prompting not only fails to consistently enhance MIA, it sometimes even underperforms compared to simple prompting. For example, GPT-4o achieves an MIA of 68.6% with CoT prompting compared to 82.8% with simple prompting for Reading. These results suggest that the reasoning steps provided by CoT prompting can be detrimental for robust counterfactual reasoning.
Our experiments reveal a decline in LLM performance with increasing DOK and grade level, highlighting challenges in handling complex questions and deep reasoning for the Malgorithm identification task. Probability emerges as the most challenging content area, while geometry is the least. This trend underscores the struggle of LLMs to sustain high performance as cognitive demands escalate.
Additionally, our experiments indicate that LLMs performance declines as the DOK and grade level increases, suggesting that the complexity of the questions and the depth of reasoning pose greater challenges for the Malgorithm identification task. Among the content areas, Probability proves to be the most difficult for LLMs, with Geometry being the least. This pattern shows that as the cognitive demands of the tasks increase, LLMs struggle more to maintain high performance levels. Overall, our findings further emphasize the need for research and development of models that can effectively engage in counterfactual reasoning.
2 Dataset: MalAlgoQA
We introduce MalAlgoQA, a real-world dataset consisting of multiple-choice questions in both mathematics and reading comprehension for students in grade 3 through 11. The dataset is designed to evaluate the counterfactual reasoning abilities of LLMs through the task of analyzing the rationales and reasons behind incorrect answer choices, a process we refer to as “Malgorithm identification". The MalAlgoQA dataset provides valuable resources to assess the counterfactual reasoning capabilities of LLMs through a pedagogical approach.
2.1 Dataset Overview
MalAlgoQA comprises two multiple-choice question sets: a mathematics and a reading comprehension question set. The mathematics question set contains 807 questions, while the reading comprehension question set has 290 questions. Each question is associated with four answer choices, as well as their corresponding rationales. Reading question set also provides a passage for each question. Table 1 provides an example of math question, with more examples available in BC. Additionally, the dataset provides rich metadata including the grade level, content classification, and Depth of Knowledge (DOK) level for each question.
2.1.1 Grade Levels
The questions in MalAlgoQA span across multiple grade levels, ranging from Grade 3 to Grade 11. As illustrated in Table 2, the distribution of questions for both mathematics and reading comprehension subsets is balanced across different grade levels.
Grade Level Math Reading
3 132 45
4 120 47
5 94 33
6 140 41
7 159 35
8 128 42
10 - 47
11 34 -
Table 2: Distribution of questions across grade levels.
2.1.2 Content Classifications
The mathematics questions set are categorized into five content classifications: Number & Operation, Algebra, Geometry & Measurement, Data Analysis, and Data Analysis & Probability. The reading comprehension questions are divided into two content classifications: Informational Text and Literature. Table 3 presents an overview of the distribution across content classifications for both question sets.
Math Content Classifications
Classification Questions
Algebra 315
Number & Operation 312
Geometry & Measurement 122
Probability 44
Data Analysis 14
Reading Content Classifications
Classification Questions
Informational Text 190
Literature 100
Table 3: Distribution of questions across content classifications.
2.1.3 Depth of Knowledge (DOK) Levels
Each question in MalAlgoQA is annotated with a DOK level, which indicates the cognitive complexity of the question. The DOK levels are derived from Norman L. Webb’s taxonomy [25] and range from I to III, with higher levels representing increased cognitive complexity. Table 4 shows the distribution of questions across DOK levels for both question sets.
DOK Level Math Reading
I 349 37
II 434 203
III 24 50
Table 4: Distribution of questions across DOK levels.
2.2 Question and Rationale Characteristics
The questions in MalAlgoQA vary in length and complexity. The average length of mathematics questions is 122 characters, while the average length of reading comprehension questions is 104 characters. Reading comprehension questions are accompanied by passages, which have an average length of 5783 characters. Each answer choice is associated with an rationale, which provide explanations for the correct and incorrect answer choices. The average length of answer rationales for mathematics questions is 227 characters, while the average length of answer rationales for reading comprehension questions is 1699 characters.
3 Methodology
In this section, we present the task formulation for Malgorithm Identification, which aims to evaluate the performance of LLMs by identifying the relationship between answer choices and their underlying rationales.
3.1 Definitions and Task Formulation
Let $Q=\{q_{1},q_{2},...,q_{n}\}$ be a set of multiple-choice questions, where each question $q_{i}$ has a set of answer choices $C_{i}=\{c_{i1},c_{i2},c_{i3},c_{i4}\}$ and a set of rationales $R_{i}=\{r_{i1},r_{i2},r_{i3},r_{i4}\}$ corresponding to each answer choice. The rationale can be thought of as a chain of reasoning steps, where each step represents a part of the thought process leading to the selection of an answer choice.
CoT Prompt for Malgorithm Identification:
System: Provide step-by-step reasoning to determine the answer rationale that corresponds to the given choice for the question. Then, provide your answer in the specified JSON format.
Question: {question}
Choice: {choice}
Answer Rationales:
{formatted_rationales}
Given the question and answer rationales, determine which answer rationale corresponds to the given choice for the question.
Provide step-by-step reasoning. Show the steps.
And then provide your answer in the following JSON format:
"Correct Choice": "[A/B/C/D]"
Table 5: Prompt for the Malgorithm Identification experiment, where the LLM identifies the relationship between answer choices and rationales in a multiple-choice question-answering setting. The task involves identifying a "malgorithm", a rationale that represents flawed reasoning leading to an incorrect answer choice. The prompt provides the question, an answer choice (either correct or incorrect), and all rationales, asking the LLM to identify the rationale corresponding to the given choice. The highlighted sentence instructs the model to provide step-by-step reasoning, an instruction added for the Chain-of-Thought prompting strategy.
Malgorithm
We introduce the term "malgorithm" to describe a rationale that contains a combination of correct reasoning steps and one or more flawed steps, which we call "mal-rules,". These "mal-rules" lead to incorrect answer choices. The term "mal-rule" is derived from the cognitive error literature [4, 23, 11], which refers to the flawed or incorrect rules that students may apply when solving problems, resulting in systematic errors. In our context, a malgorithm is a reasoning chain that consists of both correct rules and mal-rules, with the presence of mal-rules ultimately leading to the selection of an incorrect answer choice.
Formally, let $R_{i}^{correct}\subseteq R_{i}$ be the set of rationales that correspond to correct answer choices for question $q_{i}$, and let $R_{i}^{incorrect}\subseteq R_{i}$ be the set of rationales that correspond to incorrect answer choices. A rationale $r_{ij}\in R_{i}^{incorrect}$ is considered a malgorithm if it contains a combination of correct reasoning steps and one or more mal-rules that lead to the selection of an incorrect answer choice.
Malgorithm Identification Task
Given a question $q_{i}$, an answer choice $c_{ij}$ (either correct or incorrect), and the set of rationales $R_{i}$, the LLM must identify the rationale $r_{ik}$ that corresponds to the given answer choice $c_{ij}$. If the given answer choice is incorrect, the corresponding rationale is a malgorithm containing a combination of correct reasoning steps and one or more mal-rules.
3.2 Evaluation Metrics
To gain a comprehensive understanding of the LLMs performance in the Malgorithm Identification task, we use two key metrics:
1. Algorithm Identification Accuracy (AIA): This metric measures the model’s ability to identify the rationale or algorithm that aligns with a correct answer choice. Essentially, AIA evaluates how well the model recognizes and follows a chain of reasoning steps that lead to the correct answer choice.
2. Malgorithm Identification Accuracy (MIA): This metric measures the model’s ability to identify the flawed reasoning or malgorithm that aligns with an incorrect answer choice. Specifically, MIA evaluates the model’s capacity to identify faulty reasoning chains that contain a combination of correct steps and one or more mal-rules, which ultimately lead to a wrong answer.
These metrics are crucial for analyzing reasoning abilities the LLMs. AIA is essential for understanding the models’ thought processes behind correct problem-solving strategies. In contrast, MIA evaluates the model’s ability identifying flawed reasoning and counterfactual logic that lead to incorrect answers. The combination of AIA and MIA enables us to analyze the models’ performance in both causal and counterfactual reasoning.
4 Experiments
LLM Subject Num Shots Prompting MCQ AIA MIA
GPT 4o Reading 0 Simple 95.86 98.23 82.80
0 CoT 95.52 97.88 68.63
LLaMA3-70B Reading 0 Simple 89.45 95.72 60.00
0 CoT 89.27 94.60 58.40
GPT-3.5 Reading 0 Simple 84.83 95.05 13.98
0 CoT 81.66 85.82 8.61
LLaMA3-8B Reading 0 Simple 73.96 89.62 19.89
0 CoT 78.05 89.52 15.47
GPT 4o Math 0 Simple 90.11 92.06 62.82
5 Simple 90.61 64.99
0 CoT 97.61 95.65 66.10
5 CoT 95.49 65.10
LLaMA3-70B Math 0 Simple 73.83 94.31 50.22
5 Simple 93.65 48.31
0 CoT 90.72 95.47 55.32
5 CoT 95.89 50.6
GPT 3.5 Math 0 Simple 52.85 77.08 35.8
5 Simple 76.85 34.61
0 CoT 82.53 86.28 14.74
5 CoT 81.87 17.06
LLaMA3-8B Math 0 Simple 45.79 75.29 34.81
5 Simple 80.32 38.41
0 CoT 70.68 83.07 27.05
5 CoT 83.05 31.39
Table 6: Performance results of different Large Language Models (LLMs) on the Malgorithm Identification task, presented separately for Reading and Math subjects. The results are organized according to the type of prompting (Simple or CoT), and the number of shots (0 or 5). The metrics evaluated include Multiple-Choice Question (MCQ) accuracy, Algorithm Identification Accuracy (AIA), and Malgorithm Identification Accuracy (MIA). AIA measures the model’s ability to identify the correct rationale or algorithm that aligns with the given correct answer choice, whereas MIA measures the model’s ability to identify the flawed reasoning or malgorithm that aligns with the given incorrect answer choice. It can be observed that achieving higher scores in MIA is generally more challenging than AIA, highlighting the complexity of counterfactual reasoning tasks.
Subject Grade GPT 4o LLaMA3-70B
Math 3 67.00 44.26
4 65.93 54.13
5 66.52 46.97
6 67.80 58.84
7 64.50 57.49
8 62.96 42.44
11 53.68 41.03
average 65.10 50.60
Subject Grade GPT 4o LLaMA3-70B
Reading 3 73.81 65.18
4 68.89 60.71
5 69.47 45.68
6 58.62 50.00
7 63.46 80.22
8 80.67 62.00
10 64.96 46.83
average 68.63 58.40
Table 7: Grade-level performance distribution (MIA) for GPT-4o and LLaMA3-70B models using Chain-of-Thought prompting (few-shot for Math, zero-shot for Reading). Grades with above-average performance are underlined.
4.1 Experimental Setup
We conducted experiments on the Malgorithm Identification task with four state-of-the-art LLMs: GPT-4o, GPT-3.5, LLaMA3-70B and LLaMA3-8B. [^1] The specific models we use are GPT-4o: gpt-4o-2024-05-13, GPT-3.5: gpt-3.5-turbo-0125, LLaMA3-70B: Meta-LLaMA3-70B-Instruct, LLaMA3-8B: Meta-LLaMA3-8B-Instruct To evaluate the LLMs performance, we use AIA and MIA metrics (described in section 3.2).
For each model, we evaluated with four experimental settings: (1) Simple prompting, where the LLMs are prompted to provide only the final answer; (2) Chain-of-Thought (CoT) prompting, where the LLMs are prompted to provide step-by-step reasoning and show the intermediate steps before giving the final answer, as shown in Table 5; (3) Zero-shot learning, where the LLMs are provided without any additional examples; (4) Few-shot learning, where the LLMs are provided with a small number of examples demonstrating the desired reasoning process, in addition to the standard prompts. Due to context length constraints, we conducted few-shot experiments only on the Math dataset.
4.2 Effectiveness of CoT Prompting
In contrast to simple prompting strategies, Chain-of-Thought prompting requires intermediate reasoning steps that guide the model towards the final answer. CoT has demonstrated success in enhancing the reasoning capabilities of language models across various tasks [27, 7]. Nevertheless, its effectiveness in identifying flawed reasoning behind incorrect answers remain unexplored. Our objective is to determine whether the additional reasoning steps provided by CoT prompting could enhance the models ability to identify malgorithms, which is crucial for assessing models’ robustness in counterfactual reasoning and understanding student mistakes.
In Table 6, we observe that for GPT-4o, Simple prompting performs similarly to CoT prompting in detecting malgorithms for Math (65.10 vs 64.99). However, in Reading dataset, Simple prompting significantly outperforms CoT prompting in identifying malgorithms (82.80 vs 68.63). In terms of AIA, CoT prompting generally performs as well or better than Simple prompting across both subjects.. For LLaMA3-70B, Simple prompting shows better MIA than CoT prompting across both Math (50.6 vs 48.31) and Reading (60.0 vs 58.4), while AIA results remain relatively comparable.
The results demonstrates that while CoT prompting could help the identification of correct algorithms, its additional reasoning steps are not necessarily beneficial for recognizing flawed reasoning when the given answer choice is incorrect. One possible explanation is that the structured nature of CoT prompting might constrain the model’s ability to explore alternative scenarios and reason about the cause of the incorrect choice. For counterfactual reasoning, the model needs to consider possibilities that deviate from the correct reasoning path, and the intermediate steps in CoT prompting may inadvertently anchor the model to a specific line of thought, making it harder to identify the flawed reasoning behind incorrect choices. Furthermore, the models’ training data may not have sufficient examples of flawed reasoning patterns, making it challenging for the models to generate appropriate intermediate reasoning steps that lead to the identification of malgorithms, even with CoT prompting. These findings underscore the importance of malgorithm identification, which evaluates LLMs on incorrect answer choices to enhance their robustness in counterfactual reasoning.
4.3 Comparison of LLMs
Table 6 shows that malgorithm detection remains a challenging task, even for GPT-4o. This difficulty indicates room for improvement in counterfactual reasoning with LLMs. Although GPT-4o achieve over 95 in answering the multiple-choice questions in MalAlgoQA, the MIA performance drops below 70. This decline suggests that the underlying issue lies in counterfactual reasoning capabilities of the models instead of the complexity of the questions.
Furthermore, larger models such as GPT-4o and LLaMA3-70B significantly outperform smaller ones in the MIA task, while in the AIA task, performance levels are comparable across model sizes. For instance, in the Math subject with zero-shot CoT prompting, GPT-4o scored 66.10 compared to 14.74 from GPT-3.5, showcasing the large disparity. In contrast, for the AIA task, GPT-4o achieved a score of 95.65 with zero-shot CoT prompting, only marginally better than GPT-3.5’s 86.28. This suggests that the AIA task is less dependent on model size. The MIA task, which requires models to delve deeper into the underlying rationale of incorrect choices, showcases the emergent abilities of large language models [26]. This indicates that counterfactual reasoning tasks are more solvable with the enhanced capabilities of larger models.
Additionally, smaller models obtain more substantial gains with few-shot learning, especially for the MIA performance. This result suggests that for the task of malgorithm identification, smaller models could benefit more from additional context.
5 Discussions
Figure 1: An illustration of MIA performance across different Grades and DOK levels for Math and Reading question sets. The top number represents MIA performance and the bottom number represents question counts. DOK I and II are selected for Math and DOK II and III for Reading. (best viewed in colors).
Performance across Grade Levels
Table 7 analyzes the MIA performance of GPT-4o and LLaMA3-70B models across different grade levels in math and reading datasets. Both models consistently show a decline in performance with increasing grade levels, with LLaMA3-70B exhibiting some unexpected variability, particularly excelling in Grade 8 of the reading dataset. These results highlight the increasing challenges in tackling malgorithm identification task as the questions becomes more complicated.
Performance across DOK
Figure 1 presents a comprehensive overview of performance across DOK levels for various Grade using MIA score from GPT 4o with CoT prompting strategy. The pattern indicates that as the DOK level and Grade increases, MIA performance tends to decline, especially highlighted by the comparison between the top-left cell (dark green) and bottom-right cell (dark red). Similarly, in the Reading dataset, performance at DOK Level III drops to 33.33 for Grade 7. This trend further emphasizes the inherent difficulty in evaluating the malgorithms of questions that require deeper understanding and analytical thinking.
Performance across Content Classifications
Figure 2 illustrates the performance of LLMs in detecting malgorithms across five math contents and two reading content classifications. For the math dataset, Geometry consistently yields the highest performance across all models, suggesting that even for malgorithm identification, structural content like Geometry is easier for LLMs to handle as compared to numerical content. On the other hand, Probability appears to be more challenging content, reflecting the inherent difficulty of LLMs in probabilistic reasoning. For the reading dataset, GPT-4o and LLaMA3-70B excel at Informational Text, while smaller models like GPT-3.5 and LLaMA3-8B perform better with Literature. This observation suggests that larger models are better at processing structured, factual content, whereas smaller models are more efficient with narrative, context-rich material.
Figure 2: An illustration of MIA performance for LLMs across five math content classifications (Number & Operation, Algebra, Geometry & Measurement, Data Analysis, Probability) and two reading content classifications (Informational Text, Literature).
6 Related Works
Counterfactual reasoning is a cognitive process that involves considering alternative scenarios to understand causality, which has gained increasing attentions with the advent of artificial intelligence. Pearl’s foundational work on causal inference [12] has established a theoretical framework for understanding counterfactuals, which has been applied to various domains, including fairness, interpretability, and robustness in machine learning [8, 17, 24].
In the field of NLP, counterfactual reasoning is crucial for tasks that require understanding hypothetical scenarios and the causal relationships between events. However, most existing NLP models focus on predictive tasks rather than for hypothetical reasoning [3, 2, 6]. While several datasets have been created to evaluate different aspects of reasoning in language models, relatively few focus explicitly on counterfactual reasoning. For instance, the bAbI dataset [28] includes tasks that require basic forms of logical reasoning, while the SQuAD dataset [15] focuses on reading comprehension. However, these datasets do not emphasize counterfactual scenarios. Recent efforts like the COPA dataset [16] and the CQA dataset [22] have introduced questions that require commonsense reasoning, which often involves understanding causal relationships. Nonetheless, these datasets are not explicitly designed to test counterfactual reasoning. To address this gap, we introduce MalAlgoQA, a novel dataset specifically designed to evaluate the counterfactual causal reasoning capabilities of LLMs through a pedagogical approach with multiple-choice question answering format.
Applications of MalAlgoQA for personalized education: The use of LLMs in classrooms [20, 21], for instance, could profoundly influence the learning experiences of students. While LLMs can provide personalized learning support and instant feedback, their inability to accurately identify and address student misconceptions, as highlighted by our study, could potentially reinforce incorrect understanding or reasoning. However, the transparency of LLMs’ decision-making process is a crucial ethical consideration. If LLMs are to be trusted educational aids, it is critical for students, teachers, and parents to understand the reasoning behind the LLMs’ responses. Our work with the MalAlgoQA dataset is a significant step towards understanding and addressing these ethical considerations. By examining the counterfactual causal reasoning abilities of LLMs, we can assess their readiness for use in educational settings and contribute to the ongoing discourse surrounding the ethical deployment of AI in education. However, we acknowledge that this is a complex, multifaceted issue that requires continuous exploration, vigilance, and open dialogue among researchers, educators, policymakers, and society at large.
7 Conclusion and Future Works
This paper presents MalAlgoQA, a novel dataset designed to critically evaluate the counterfactual reasoning capabilities of LLMs through a pedagogical prespective. Additionally, we introduce the Malgorithm Identification task along with associated metrics for its evaluation. Our findings underscore the need for ongoing research and development the ability to identify malgorithms, particularly for educational applications where recognizing flawed reasoning processes is essential. Notably, our study also challenges the efficacy of the chain-of-thought prompting technique in enhancing counterfactual causal reasoning, advocating for further investigation into suitable prompting techniques for robust malgorithm identification. In the future, we plan to expand the MalAlgoQA dataset to cover a broader range of subjects and reasoning complexities. We also aim to explore different training methodologies and architectural modifications to LLMs that may enhance their ability to engage in counterfactual causal reasoning, ultimately contributing to the development of more cognitively-aligned language models.
Acknowledgments
This work was supported by NSF grants 1842378, ONR grant N0014-20-1-2534, AFOSR grant FA9550-22-1-0060, and a Vannevar Bush Faculty Fellowship, ONR grant N00014-18-1-2047.
Appendix A Additional Analysis
A.1 Grade vs DOK for different content classifications
Figure 3: An image of MIA performance across different Grades and DOK levels for various Content Classifications for Math question sets. The top number represents MIA performance and the bottom number represents question counts. The data analysis and probability classifications are combined together for illustration. (best viewed in colors).
Figure 4: An image of MIA performance across different Grades and DOK levels for two Content Classifications for Reading question sets. The top number represents MIA performance and the bottom number represents question counts (best viewed in colors).
Figures 3 and 4 illustrate the relationships between grade levels, DOK for different content classifications in math and reading question sets. In mathematics, Number & Operation and Algebra questions demonstrate relatively consistent performance, with higher DOK levels often showing better results. For instance, MIA performance for Number & Operation peaks in lower grades and declines in higher grades, particularly Grade 7. This trend suggests increasing complexity in numerical operations as grade levels increases. On the other hand, Geometry & Measurements, Data Analysis and Probability exhibit high variability, with strong MIA performance in some grades but gaps in others. In reading, performance for Informational Text is generally consistent across grades, peaking in Grade 8, with lower DOK levels typically yielding higher MIA performance. Literature shows high variability, reflecting the varying complexity of literary analysis tasks. Notably, content classifications in mathematics show more consistent patterns across grades compared to reading ones.
A.2 Performance across Grade level and Content Classifications
Figure 5: An illustration of MIA performance across different Grades and Content Classifications for Math and Reading question sets. The top number represents MIA performance and the bottom number represents question counts (best viewed in colors).
Figure 5 shows the relationship between grade levels and content classifications for math and reading question sets. In mathematics question set, Algebra demonstrates relatively consistent performance across all grades, with the number of algebra questions increasing significantly in higher grades, peaking at 191 data points in Grade 8. Number & Operation exhibits a gradual decline in MIA performance from lower to higher grades. In contrast, Data Analysis and Probability, Geometry & Measurement classifications show varied MIA performance across grade levels. For instance, Data Analysis and Probability demonstrates exceptionally high performance in Grade 6 (81.82), albeit with relatively few data points (22). In reading question set, MIA performance for both Informational Text and Literature classifications fluctuates across grades without much correlation. Informational Text achieves a significant peak in MIA performance for Grade 8, while Literature questions in Grade 5 demonstrate exceptionally high performance. This variability could be attributed to differences in text complexity or question difficulty at different grade levels for the Malgorithm Identification problem.
A.3 Performance across Questions
Figure 6: The distribution of GPT4-o performance in identifying malgorithms across question levels for math and reading question sets. The x-axis represents the number of correctly identified malgorithms per question, ranging from 0 to 3.
We can find the performance of LLMs in identifying malgorithms across each question in both mathematics and reading comprehension sets in Figure 6. The number of questions increases as we progress from 0 to 3 correctly identified malgorithms per question. This pattern suggests that LLMs are more likely to identify multiple malgorithms correctly rather than missing all or most of them. However, despite this positive trend, it is crucial to note that LLMs’ overall performance on this task remains suboptimal. The presence of the significant number of questions at the 0 and 1 levels across both domains indicates room for improvement. In mathematics, the combined number of questions with 0 or 1 correctly identified malgorithms is comparable to those with 3 correct identifications. Similarly, in reading comprehension, while there is a clearer upward trend, the number of questions with 0 or 1 correct identifications remains considerable. These findings underscore the need for further research of LLMs’ counterfactual reasoning capabilities in malgorithm identification across various question types and subject areas.
Appendix B Examples of Mathematics dataset
We provide five examples, one per content classifications, of question, choices and their corresponding rationales in Table 8 and 9.
Math Question Set Example 1:
Content Classification: Number & Operation
Question: What number is subtracted from 1,000 to result in a difference of 421?
Answer and Rationales:
A: 421
Rationale A: Selects the result of subtraction.
B: 579
Rationale B: 1000 - 421 = 579
C: 621
Rationale C: Rounded the hundreds place and added the tens and ones place. 1000 - 400 = 600; 600 + 21 = 621
D: 689
Rationale D: Rounded in hundreds, tens and ones place. 1000 - 400 = 600; 100 - 20 = 80; 10 - 1 = 9; 600 + 80 + 9 = 689
Math Question Set Example 2:
Content Classification: Algebra
Question: Anne bought a calculator that cost 30. She received 10% off her purchase and then was charged 6% tax. What was the total amount that Anne paid?
Answer and Rationales:
A: 21.20
Rationale A: Subtracted 10 from 30, then added 6%.
B: 25.38
Rationale B: Subtracted 10% of the price from 30 then subtracted 6%.
C: 28.62
Rationale C: Subtracted 10% of the price from 30 and then added 6%.
D: 28.80
Rationale D: Calculated tax based on original price, then subtracted the 3.
Math Question Set Example 3:
Content Classification: Geometry and Measurement
Question: Which side lengths form a right triangle?
Answer and Rationales:
A: 2 cm, 4 cm, 8 cm
Rationale A: Multiplied the 2 smaller sides to get the longest side.
B: 4 cm, 5 cm, 6 cm
Rationale B: Selected side with constant difference between the sides.
C: 5 cm, 12 cm, 13 cm
Rationale C: 5^2 + 12^2 = 13^2
D: 9 cm, 16 cm, 25 cm
Rationale D: Added the 2 smaller sides to get the longest side.
Table 8: Examples of Mathematics question set.
Math Question Set Example 4:
Content Classification: Data Analysis
Question: Mrs. Castillo recorded the number of students in each grade 5 classroom.
24 28 25 24 29
What is the mean of the data?
Answer and Rationales:
A: 24
Rationale A: Chose the mode.
B: 25
Rationale B: Chose the median. {24, 24, 25, 28, 29}
C: 26
Rationale C: (24 + 28 + 25 + 24 + 29) / 5 = 130/5 = 26
D: 29
Rationale D: Chose the greatest number.
Math Question Set Example 5:
Content Classification: Probability
Question: Ms. Collier had a deck of cards. There were stars on 1/4 of the cards in the deck. After randomly picking a card and returning it to the deck 100 times, the expected result and relative frequency of picking a card with a star were equal. In the first 50 cards she picked, she got a star 10 times. How many times did Ms. Collier get a star in the second 50 cards she picked?
Answer and Rationales:
A: 10
Rationale A: Chose the number of times a star is picked in the 1st 50 cards drawn.
B: 13
Rationale B: Calculated 50 * 1/4 = 12.5 and then rounded to 13.
C: 15
Rationale C: Correct. 100 * 1/4 = 25 cards w/ stars. 25 - 10 = 15
D: 25
Rationale D: Calculated 100 * 1/4 = 25 cards w/ stars.
Table 9: Examples of Mathematics question set.
Appendix C Examples of Reading Comprehension dataset
We provide two examples of passage, question, choices and their corresponding rationales in Table 10 and 11.
Reading Question Set Example 1:
Content Classification: Literature
Passage: Blueberries
1 You ought to have seen what I saw on my way
2 To the village, through Pattersons pasture to-day:
3 Blueberries as big as the end of your thumb,
4 Real sky-blue, and heavy, and ready to drum
5 In the cavernous pail of the first one to come!
6 And all ripe together, not some of them green
7 And some of them ripe! You ought to have seen!
Question: In lines 1-7, the speaker expresses great surprise about the ...?
Answer and Rationales:
A: large pails of blueberries.
Rationale A: Some readers might find this option plausible because the speaker mentions the cavernous pail of the first one to come" (line 5). However, this mention of the pail is referring to the pail into which the lucky picker who first stumbles across this crop will put the Blueberries as big as the end of your thumb" (line 3). The speaker does not express surprise that pickers are bringing large pails.
B: number of blueberry pickers.
Rationale B: Some readers might choose this option because the phrase the first one to come" in line 5 refers to a blueberry picker. However, although the speaker is predicting success for anyone who comes to pick the blueberries, the speaker is not expressing surprise that there are no pickers yet. Based on the speakers description of how big, blue, and ripe the berries are (as big as the end of your thumb, Real sky-blue, and heavy, and ready" (lines 3 and 4)), the berries are undisturbed, and no one has discovered them except for the speaker.
C: ripeness of the blueberries.
Rationale C: The speaker describes big as the end of your thumb" (line 3) blueberries that are Real sky-blue, and heavy," (line 4) and all ripe together" (line 6). The speaker repeats, You ought to have seen" in lines 1 and 7 to convey surprise.
D: size of new blueberry bushes.
Rationale D: This option might seem plausible because the size of the blueberries and their bushes are described. The first speaker draws attention to size when describing the blueberries as big as the end of your thumb" (line 3). In line 13, the second speaker exclaims, Why, there hasnt been time for the bushes to grow." This line implies that the bushes must be small, so it is surprising that the blueberries they produce are numerous and as big as the end of your thumb." Although lines 1-7 address the blueberries, the speaker does not mention the bushes on which they grow in those lines.
Table 10: Example 1 of Reading question set.
Reading Question Set Example 2:
Content Classification: Informational Text
Passage: Creating Codes: Grace Hopper and the Computer Compiler
1 The computer screen freezes. No matter what or how often you click, the hourglass or rotating circle (that symbol that insists, Im working on it!") spins away. You want to shake the screen. Someone next to you says, I think you found a bug."
Question: What is the main purpose of paragraph 1?
Answer and Rationales:
A: To demonstrate the importance of the subject
Rationale A: Although some readers may conclude that the frozen computer screen in paragraph 1 demonstrates how important computer programming is currently, the paragraph describes a familiar feeling of frustration rather than detailing the more serious consequences that could result from programming mistakes.
B: To help readers make a connection with the topic
Rationale B: The paragraph describes a person facing a frozen computer screen that many readers have probably experienced in order to help readers make a personal connection to the topic. By describing this common situation, the readers interest is engaged. So, when the article continues on to discuss Hoppers contribution to programming and how she helped solve some of the early computer bugs, readers feel more connected with the topic.
C: To give background information about the subject
Rationale C: Although introductions often provide background information, and some readers may conclude that explaining a computer bug is background information for the article, the topic of the article focuses not on computer bugs but on Hoppers contributions to computer programming.
D: To provide readers with a brief summary of the topic
Rationale D: Although the main topic of the article is Hoppers contribution to computer programming and paragraph 1 mentions computers, paragraph 1 describes a common situation in order to raise interest among readers. It describes a person having computer problems rather than summarizing Hoppers contributions to computer programming.
Table 11: Example 2 of Reading question set

### v2

@@
-MalAlgoQA: A Pedagogical Approach for Evaluating Counterfactual Reasoning Abilities
+MalAlgoQA: Pedagogical Evaluation of Counterfactual Reasoning in Large Language Models and Implications for AI in Education
 Abstract
@@
 The dataset comprises mathematics and reading comprehension questions, each accompanied by four answer choices and their corresponding rationales.
-We focus on the incorrect answer rationales, termed "malgorithms", which highlights flawed reasoning steps leading to incorrect answers and offers valuable insights into erroneous thought processes.
-We also propose the Malgorithm Identification task, where LLMs are assessed based on their ability to identify corresponding malgorithm given an incorrect answer choice.
+At the heart of MalAlgoQA are “malgorithms” - rationales behind incorrect answer choices that represent flawed yet logically coherent reasoning paths.
+These malgorithms serve as counterfactual scenarios, allowing us to assess an LLM’s ability to identify and analyze flawed reasoning patterns.
+We propose the Malgorithm Identification task, where LLMs are assessed based on their ability to identify corresponding malgorithm given an incorrect answer choice.
 To evaluate the model performance, we introduce two metrics: Algorithm Identification Accuracy (AIA) for correct answer rationale identification, and Malgorithm Identification Accuracy (MIA) for incorrect answer rationale identification.
-The task is challenging since state-of-the-art LLMs exhibit significant drops in MIA as compared to AIA.
-Moreover, we find that the chain-of-thought prompting technique not only fails to consistently enhance MIA, but can also lead to underperformance compared to simple prompting.
-These findings hold significant implications for the development of more cognitively-inspired LLMs to improve their counterfactual reasoning abilities, particularly through a pedagogical perspective where understanding and rectifying student misconceptions are crucial.
-The data and code can be found at https://github.com/luffycodes/MalAlgoQA-Dataset.
+Our experiments reveal that state-of-the-art LLMs exhibit significant performance drops in MIA compared to AIA, highlighting the challenges in counterfactual reasoning.
+Surprisingly, we find that the chain-of-thought prompting technique not only fails to consistently enhance MIA but can sometimes lead to underperformance compared to simple prompting.
+These findings have important implications for developing LLMs with improved counterfactual reasoning, particularly relevant for AI-powered tutoring systems, where identifying and addressing student misconceptions is essential.
+MalAlgoQA dataset is available here.
+MalAlgoQA: Pedagogical Evaluation of Counterfactual Reasoning in
+Large Language Models and Implications for AI in Education
+Naiming Liu Rice University nl35@rice.edu                        Shashank Sonkar Rice University ss164@rice.edu
+MyCo Le Rice University msl12@rice.edu                        Richard G. Baraniuk Rice University richb@rice.edu
+footnotetext: Equal contribution.
 1 Introduction
-Counterfactual reasoning refers to the ability to consider alternative scenarios and infer the cause of an event based on hypothetical situations.
-It is a fundamental aspect of human cognition that plays a crucial role in problem-solving, decision-making, and learning [5, 13, 14].
-Although large language models (LLMs) have demonstrated remarkable performance in various natural language processing tasks, their ability to engage in counterfactual reasoning remains relatively underexplored.
-Counterfactual reasoning requires a degree of hypothetical and abstract thinking, which poses a significant challenge for current LLMs, as they are generally trained to predict outcomes based on facts, rather than imagining alternative scenarios that might not be explicitly present in the training data [3, 9, 19].
+Evaluating the counterfactual reasoning capabilities of Large Language Models (LLMs) – their ability to consider “what if” scenarios and infer causes based on hypothetical situations – remains a significant challenge despite their impressive performance in various natural language tasks (Pearl, 2019; Rafetseder et al., 2021; Sonkar and Baraniuk, 2023).
+We introduce a novel method to evaluate these capabilities, drawing inspiration from educational assessment techniques.
+Specifically, we leverage the practice of distractor choice generation, which involves creating plausible but incorrect answer options by envisioning hypothetical scenarios and logical yet flawed reasoning paths.
+This approach mimics how educators anticipate and model potential student misconceptions, engaging in counterfactual thinking about alternative cognitive processes that could lead to specific errors.
 Question Each side of a square is 3 inches long. A student increases the length of each side by 3 inches. The area of the new square is how many times the area of the original square?
@@
 For incorrect answer choices, the corresponding rationales are malgorithms, representing flawed reasoning processes and thereby evaluating the counterfactual reasoning abilities of LLMs.
-To fill this gap, we introduce MalAlgoQA, a novel dataset designed to evaluate the counterfactual reasoning capabilities of LLMs through a pedagogical approach with multiple-choice question answering.
-MalAlgoQA consists of a diverse set of 807 mathematics and 290 reading comprehension questions spanning various grade levels (3-11), content classifications (e.g., algebra, geometry, number & operations), and Depth of Knowledge levels (DOK) (1-3) [25].
-Each question is accompanied by a set of answer choices and their associated rationales, which we term "malgorithms" (inspired by the concept of "mal-rules" proposed by [11]).
-Malgorithms represent the flawed reasoning that leads to an incorrect answer, revealing the underlying thought process that causes a student to arrive at the incorrect conclusion.
-With MalAlgoQA, we propose the Malgorithm Identification task, which involves identifying the rationale behind the selection of a particular answer choice.
-When the given answer choice is correct, a model needs to employ causal reasoning to determine the rationale or algorithm that justified the correct selection.
-On the other hand, when the given answer choice is incorrect, the task demands counterfactual reasoning, as the model must identify the flawed reasoning or malgorithm that led to the selection of an incorrect option.
-In this case, the model must consider hypothetical scenarios and reason about the cause of an incorrect choice, which may not align with the facts.
-To illustrate the Malgorithm Identification task, consider a simple question: "What is 1 + 2 × 3 + 4?" with choices A.
+Building on this foundation, we present MalAlgoQA, a dataset designed to challenge LLMs’ ability to reason about hypothetical situations and identify flawed logical pathways.
+MalAlgoQA consists of a diverse set of 807 mathematics and 290 reading comprehension questions spanning various grade levels (3-11), content classifications (e.g., algebra, geometry, number & operations), and Depth of Knowledge levels (DOK) (1-3) (Webb, 2002).
+Each question is accompanied by a set of answer choices and their associated rationales.
+Central to our approach is the concept of “malgorithms” - a term we coin to describe the flawed reasoning paths that lead to incorrect answers (name is inspired by the concept of “mal-rules” proposed by Payne and Squibb (1990)).
+These malgorithms represent deterministic yet erroneous thought processes, mirroring the way students might logically follow a series of steps based on a misunderstanding or misapplication of concepts.
+We propose the Malgorithm Identification task to assess LLMs’ ability to reason about both correct and incorrect answer rationales.
+In this task, given a question and a specific answer choice, the model must identify the underlying rationale (algorithm or malgorithm) that led to that choice.
+This approach challenges LLMs to engage in both causal and counterfactual reasoning.
+To illustrate, consider a simple question: “What is 1 + 2 × 3 + 4?” with choices A.
 11, B.
@@
 21.
-The rationales are: "Worked left to right" (corresponding to choice B), "Applied PEASMD rule instead of PEMDAS" (corresponding to choice C), and "Applied PEMDAS rule" (corresponding to the correct choice A).
+The rationales are: “Worked left to right” (corresponding to choice B), “Applied PEASMD rule instead of PEMDAS” (corresponding to choice C), and “Applied PEMDAS rule” (corresponding to the correct choice A).
 Given an answer choice, the model should identify the rationale explaining that choice.
@@
 Malgorithm Identification Accuracy (MIA), of particular interest, evaluates the model’s counterfactual causal reasoning abilities by quantifying accuracy in identifying the rationale behind an incorrect answer.
-Our experiments reveal that GPT-4o [10], a state-of-the-art LLM, achieves an AIA of $95.7\%$, demonstrating its proficiency in this task.
+Our experiments reveal that GPT-4o (OpenAI, 2023) achieves an AIA of $95.7\%$, demonstrating its proficiency in this task.
 However, when it comes to MIA, GPT-4o’s performance drops significantly, with an accuracy of only $66.1\%$.
-This pattern, where accuracy is higher for correct answer rationales and lower for incorrect answer rationales, is consistent across other language models, such as LLaMA-3-70B [1] and GPT-3.5 [18].
+This pattern, where accuracy is higher for correct answer rationales and lower for incorrect answer rationales, is consistent across other language models, such as LLaMA-3-70B (Dubey et al., 2024) and GPT-3.5 (Schulman et al., 2022).
 It emphasizes the challenge of counterfactual reasoning for LLMs, especially on finding the underlying rationale behind incorrect answer choices.
-Our experiments also reveal an unexpected result that Chain-of-Thought (CoT) prompting not only fails to consistently enhance MIA, it sometimes even underperforms compared to simple prompting.
-For example, GPT-4o achieves an MIA of 68.6% with CoT prompting compared to 82.8% with simple prompting for Reading.
-These results suggest that the reasoning steps provided by CoT prompting can be detrimental for robust counterfactual reasoning.
-Our experiments reveal a decline in LLM performance with increasing DOK and grade level, highlighting challenges in handling complex questions and deep reasoning for the Malgorithm identification task.
-Probability emerges as the most challenging content area, while geometry is the least.
-This trend underscores the struggle of LLMs to sustain high performance as cognitive demands escalate.
-Additionally, our experiments indicate that LLMs performance declines as the DOK and grade level increases, suggesting that the complexity of the questions and the depth of reasoning pose greater challenges for the Malgorithm identification task.
-Among the content areas, Probability proves to be the most difficult for LLMs, with Geometry being the least.
-This pattern shows that as the cognitive demands of the tasks increase, LLMs struggle more to maintain high performance levels.
-Overall, our findings further emphasize the need for research and development of models that can effectively engage in counterfactual reasoning.
-2 Dataset: MalAlgoQA
-We introduce MalAlgoQA, a real-world dataset consisting of multiple-choice questions in both mathematics and reading comprehension for students in grade 3 through 11.
-The dataset is designed to evaluate the counterfactual reasoning abilities of LLMs through the task of analyzing the rationales and reasons behind incorrect answer choices, a process we refer to as “Malgorithm identification".
-The MalAlgoQA dataset provides valuable resources to assess the counterfactual reasoning capabilities of LLMs through a pedagogical approach.
+Our experiments also reveal an unexpected result that Chain-of-Thought (CoT) prompting Wei et al.
+(2022b); Kojima et al.
+(2022) not only fails to consistently enhance MIA, it sometimes even underperforms compared to simple prompting.
+The implications of our findings reveal critical challenges for AI in education.
+The stark performance gap between AIA and MIA across all models (e.g., GPT-4o’s $95.7\%$ AIA vs.
+$66.1\%$ MIA) exposes fundamental limitations in LLMs’ ability to understand student misconceptions.
+This discrepancy suggests that current LLM tutoring systems fail to identify and address the root causes of student errors.
+Such limitations necessitate a rethinking of feedback mechanisms in LLM-powered educational tools.
+Their struggle with malgorithm identification (e.g., GPT-3.5’s dramatic drop from $86.3\%$ AIA to $14.74\%$ MIA in Math) indicates they may offer incomplete or potentially misleading guidance when addressing student mistakes.
+These insights have profound implications for automated grading and assessment.
+The significant performance drop in MIA across all models raises concerns about the reliability of AI-driven systems in evaluating open-ended questions or complex problem-solving tasks.
+Our work makes two key contributions to the field of LLMs and AI for education, with critical implications for developing more effective AI-powered tutoring systems.
+First, we introduce MalAlgoQA in section 2, a novel dataset inspired by pedagogical distractor generation techniques, specifically designed to evaluate the counterfactual reasoning capabilities of LLMs.
+This approach mimics how educators create plausible but incorrect answer options, challenging LLMs to engage in the type of reasoning essential for understanding and addressing student misconceptions.
+Second, we propose the Malgorithm Identification task in section 3, along with two associated evaluation metrics: AIA and MIA.
+Our findings in sections 4 and 5 reveal a significant performance gap between these metrics across various LLMs, highlighting the challenge these models face in counterfactual reasoning.
+We discuss the implications of our findings for AI-powered educational tool in section 6.
+By developing methods to evaluate LLMs’ ability to identify and understand student misconceptions, our work paves the way for more effective AI tutoring systems that can accurately diagnose and address individual students’ conceptual gaps Liu et al.
+(2023); Sonkar et al.
+(2023, 2024c).
+2 MalAlgoQA Dataset
+We introduce MalAlgoQA, a real-world dataset consisting of multiple-choice questions in mathematics and reading comprehension for students in grade 3 through 11.
+The dataset is designed to evaluate the counterfactual reasoning abilities of LLMs through the task of analyzing the rationales and reasons behind incorrect answer choices, a process we refer to as “Malgorithm identification.” The MalAlgoQA dataset provides valuable resources to assess the counterfactual reasoning capabilities of LLMs through a pedagogical approach.
 2.1 Dataset Overview
@@
 Reading question set also provides a passage for each question.
-Table 1 provides an example of math question, with more examples available in BC.
+Table 1 provides an example of math question, with more examples available in A.
 Additionally, the dataset provides rich metadata including the grade level, content classification, and Depth of Knowledge (DOK) level for each question.
@@
 The DOK levels are derived from Norman L.
-Webb’s taxonomy [25] and range from I to III, with higher levels representing increased cognitive complexity.
+Webb’s taxonomy Webb (2002) and range from I to III, with higher levels representing increased cognitive complexity.
 Table 4 shows the distribution of questions across DOK levels for both question sets.
@@
 3 Methodology
-In this section, we present the task formulation for Malgorithm Identification, which aims to evaluate the performance of LLMs by identifying the relationship between answer choices and their underlying rationales.
-3.1 Definitions and Task Formulation
-Let $Q=\{q_{1},q_{2},...,q_{n}\}$ be a set of multiple-choice questions, where each question $q_{i}$ has a set of answer choices $C_{i}=\{c_{i1},c_{i2},c_{i3},c_{i4}\}$ and a set of rationales $R_{i}=\{r_{i1},r_{i2},r_{i3},r_{i4}\}$ corresponding to each answer choice.
-The rationale can be thought of as a chain of reasoning steps, where each step represents a part of the thought process leading to the selection of an answer choice.
 CoT Prompt for Malgorithm Identification:
@@
 Choice: {choice}
-Answer Rationales:
-{formatted_rationales}
+Answer Rationales: {formatted_rationales}
 Given the question and answer rationales, determine which answer rationale corresponds to the given choice for the question.
@@
 Table 5: Prompt for the Malgorithm Identification experiment, where the LLM identifies the relationship between answer choices and rationales in a multiple-choice question-answering setting.
-The task involves identifying a "malgorithm", a rationale that represents flawed reasoning leading to an incorrect answer choice.
+The task involves identifying the “malgorithm”, a rationale that represents flawed reasoning steps which lead to an incorrect answer choice.
 The prompt provides the question, an answer choice (either correct or incorrect), and all rationales, asking the LLM to identify the rationale corresponding to the given choice.
 The highlighted sentence instructs the model to provide step-by-step reasoning, an instruction added for the Chain-of-Thought prompting strategy.
+In this section, we present the task formulation for Malgorithm Identification.
+3.1 Definitions and Task Formulation
+Let $Q=\{q_{1},q_{2},...,q_{n}\}$ be a set of multiple-choice questions, where each question $q_{i}$ has a set of answer choices $C_{i}=\{c_{i1},c_{i2},c_{i3},c_{i4}\}$ and a set of rationales $R_{i}=\{r_{i1},r_{i2},r_{i3},r_{i4}\}$ corresponding to each answer choice.
+The rationale can be thought of as a chain of reasoning steps, where each step represents a part of the thought process leading to the selection of an answer choice.
 Malgorithm
-We introduce the term "malgorithm" to describe a rationale that contains a combination of correct reasoning steps and one or more flawed steps, which we call "mal-rules,".
-These "mal-rules" lead to incorrect answer choices.
-The term "mal-rule" is derived from the cognitive error literature [4, 23, 11], which refers to the flawed or incorrect rules that students may apply when solving problems, resulting in systematic errors.
+We introduce the term “malgorithm” to describe a rationale underlying an incorrect answer choice.
+A malgorithm is composed of a series of reasoning steps, some of which are correct, while others contain errors.
+We refer to these erroneous steps as “mal-rules”.
+The term mal-rules is derived from the cognitive error literature (Brown and VanLehn, 1980; VanLehn, 1990; Payne and Squibb, 1990), which refers to the flawed or incorrect rules that students may apply when solving problems, resulting in systematic errors.
 In our context, a malgorithm is a reasoning chain that consists of both correct rules and mal-rules, with the presence of mal-rules ultimately leading to the selection of an incorrect answer choice.
-Formally, let $R_{i}^{correct}\subseteq R_{i}$ be the set of rationales that correspond to correct answer choices for question $q_{i}$, and let $R_{i}^{incorrect}\subseteq R_{i}$ be the set of rationales that correspond to incorrect answer choices.
-A rationale $r_{ij}\in R_{i}^{incorrect}$ is considered a malgorithm if it contains a combination of correct reasoning steps and one or more mal-rules that lead to the selection of an incorrect answer choice.
 Malgorithm Identification Task
@@
 1.
-Algorithm Identification Accuracy (AIA): This metric measures the model’s ability to identify the rationale or algorithm that aligns with a correct answer choice.
-Essentially, AIA evaluates how well the model recognizes and follows a chain of reasoning steps that lead to the correct answer choice.
+Algorithm Identification Accuracy (AIA): This metric measures the model’s ability to identify the rationale or algorithm that corresponds to the correct answer choice.
+Essentially, AIA evaluates how well the model recognizes and follows a chain of reasoning steps that lead to the correct answer.
 2.
-Malgorithm Identification Accuracy (MIA): This metric measures the model’s ability to identify the flawed reasoning or malgorithm that aligns with an incorrect answer choice.
-Specifically, MIA evaluates the model’s capacity to identify faulty reasoning chains that contain a combination of correct steps and one or more mal-rules, which ultimately lead to a wrong answer.
-These metrics are crucial for analyzing reasoning abilities the LLMs.
-AIA is essential for understanding the models’ thought processes behind correct problem-solving strategies.
-In contrast, MIA evaluates the model’s ability identifying flawed reasoning and counterfactual logic that lead to incorrect answers.
-The combination of AIA and MIA enables us to analyze the models’ performance in both causal and counterfactual reasoning.
-4 Experiments
+Malgorithm Identification Accuracy (MIA): This metric measures the model’s ability to identify the flawed reasoning or malgorithm that corresponds to an incorrect answer choice.
+Specifically, MIA evaluates the model’s capacity to identify faulty reasoning chains that contain a combination of correct steps and one or more mal-rules, which lead to the wrong answer.
+4 Experiments and Findings
 LLM Subject Num Shots Prompting MCQ AIA MIA
@@
 5 CoT 83.05 31.39
-Table 6: Performance results of different Large Language Models (LLMs) on the Malgorithm Identification task, presented separately for Reading and Math subjects.
-The results are organized according to the type of prompting (Simple or CoT), and the number of shots (0 or 5).
-The metrics evaluated include Multiple-Choice Question (MCQ) accuracy, Algorithm Identification Accuracy (AIA), and Malgorithm Identification Accuracy (MIA).
-AIA measures the model’s ability to identify the correct rationale or algorithm that aligns with the given correct answer choice, whereas MIA measures the model’s ability to identify the flawed reasoning or malgorithm that aligns with the given incorrect answer choice.
-It can be observed that achieving higher scores in MIA is generally more challenging than AIA, highlighting the complexity of counterfactual reasoning tasks.
+Table 6: Performance results of LLMs on the Malgorithm Identification task for Reading and Math subjects.
+Results are organized by prompting type (Simple/CoT) and number of shots (0/5).
+Metrics include Multiple-Choice Question (MCQ) accuracy, Algorithm Identification Accuracy (AIA), and Malgorithm Identification Accuracy (MIA).
+AIA measures correct rationale identification, while MIA measures flawed reasoning identification.
+MIA scores are generally significantly lower than AIA, highlighting the challenge of counterfactual reasoning.
 Subject Grade GPT 4o LLaMA3-70B
@@
 4.1 Experimental Setup
-We conducted experiments on the Malgorithm Identification task with four state-of-the-art LLMs: GPT-4o, GPT-3.5, LLaMA3-70B and LLaMA3-8B.
-[^1] The specific models we use are GPT-4o: gpt-4o-2024-05-13, GPT-3.5: gpt-3.5-turbo-0125, LLaMA3-70B: Meta-LLaMA3-70B-Instruct, LLaMA3-8B: Meta-LLaMA3-8B-Instruct To evaluate the LLMs performance, we use AIA and MIA metrics (described in section 3.2).
-For each model, we evaluated with four experimental settings: (1) Simple prompting, where the LLMs are prompted to provide only the final answer; (2) Chain-of-Thought (CoT) prompting, where the LLMs are prompted to provide step-by-step reasoning and show the intermediate steps before giving the final answer, as shown in Table 5; (3) Zero-shot learning, where the LLMs are provided without any additional examples; (4) Few-shot learning, where the LLMs are provided with a small number of examples demonstrating the desired reasoning process, in addition to the standard prompts.
+We conducted experiments on the Malgorithm Identification task with four state-of-the-art LLMs: GPT-4o, GPT-3.5, LLaMA3-70B and LLaMA3-8B.The specific models we use are GPT-4o: gpt-4o-2024-05-13, GPT-3.5: gpt-3.5-turbo-0125, LLaMA3-70B: Meta-LLaMA3-70B-Instruct, LLaMA3-8B: Meta-LLaMA3-8B-Instruct To evaluate the LLMs performance, we use AIA and MIA metrics (described in section 3.2).
+For each model, we evaluated with four experimental settings: (1) Simple prompting, where the LLMs are prompted to provide only the final answer; (2) Chain-of-Thought (CoT) prompting Wei et al.
+(2022b); Kojima et al.
+(2022), where the LLMs are prompted to provide step-by-step reasoning and show the intermediate steps before giving the final answer, as shown in Table 5; (3) Zero-shot learning, where the LLMs are provided without any additional examples; (4) Few-shot learning, where the LLMs are provided with a small number of examples demonstrating the desired reasoning process, in addition to the standard prompts.
 Due to context length constraints, we conducted few-shot experiments only on the Math dataset.
-4.2 Effectiveness of CoT Prompting
-In contrast to simple prompting strategies, Chain-of-Thought prompting requires intermediate reasoning steps that guide the model towards the final answer.
-CoT has demonstrated success in enhancing the reasoning capabilities of language models across various tasks [27, 7].
-Nevertheless, its effectiveness in identifying flawed reasoning behind incorrect answers remain unexplored.
-Our objective is to determine whether the additional reasoning steps provided by CoT prompting could enhance the models ability to identify malgorithms, which is crucial for assessing models’ robustness in counterfactual reasoning and understanding student mistakes.
-In Table 6, we observe that for GPT-4o, Simple prompting performs similarly to CoT prompting in detecting malgorithms for Math (65.10 vs 64.99).
-However, in Reading dataset, Simple prompting significantly outperforms CoT prompting in identifying malgorithms (82.80 vs 68.63).
-In terms of AIA, CoT prompting generally performs as well or better than Simple prompting across both subjects..
-For LLaMA3-70B, Simple prompting shows better MIA than CoT prompting across both Math (50.6 vs 48.31) and Reading (60.0 vs 58.4), while AIA results remain relatively comparable.
-The results demonstrates that while CoT prompting could help the identification of correct algorithms, its additional reasoning steps are not necessarily beneficial for recognizing flawed reasoning when the given answer choice is incorrect.
-One possible explanation is that the structured nature of CoT prompting might constrain the model’s ability to explore alternative scenarios and reason about the cause of the incorrect choice.
-For counterfactual reasoning, the model needs to consider possibilities that deviate from the correct reasoning path, and the intermediate steps in CoT prompting may inadvertently anchor the model to a specific line of thought, making it harder to identify the flawed reasoning behind incorrect choices.
-Furthermore, the models’ training data may not have sufficient examples of flawed reasoning patterns, making it challenging for the models to generate appropriate intermediate reasoning steps that lead to the identification of malgorithms, even with CoT prompting.
-These findings underscore the importance of malgorithm identification, which evaluates LLMs on incorrect answer choices to enhance their robustness in counterfactual reasoning.
-4.3 Comparison of LLMs
-Table 6 shows that malgorithm detection remains a challenging task, even for GPT-4o.
-This difficulty indicates room for improvement in counterfactual reasoning with LLMs.
-Although GPT-4o achieve over 95 in answering the multiple-choice questions in MalAlgoQA, the MIA performance drops below 70.
-This decline suggests that the underlying issue lies in counterfactual reasoning capabilities of the models instead of the complexity of the questions.
-Furthermore, larger models such as GPT-4o and LLaMA3-70B significantly outperform smaller ones in the MIA task, while in the AIA task, performance levels are comparable across model sizes.
-For instance, in the Math subject with zero-shot CoT prompting, GPT-4o scored 66.10 compared to 14.74 from GPT-3.5, showcasing the large disparity.
-In contrast, for the AIA task, GPT-4o achieved a score of 95.65 with zero-shot CoT prompting, only marginally better than GPT-3.5’s 86.28.
-This suggests that the AIA task is less dependent on model size.
-The MIA task, which requires models to delve deeper into the underlying rationale of incorrect choices, showcases the emergent abilities of large language models [26].
-This indicates that counterfactual reasoning tasks are more solvable with the enhanced capabilities of larger models.
-Additionally, smaller models obtain more substantial gains with few-shot learning, especially for the MIA performance.
-This result suggests that for the task of malgorithm identification, smaller models could benefit more from additional context.
-5 Discussions
+4.2 Main Findings
+Wide Performance Gap Between MIA and AIA
+The most striking observation from Table 6 is the substantial performance gap between Malgorithm Identification Accuracy (MIA) and Algorithm Identification Accuracy (AIA) across all models.
+For GPT-4o on the Math dataset with zero-shot CoT prompting, we see an AIA of $95.65\%$ compared to an MIA of only $66.10\%$ a drop of nearly $30$ percentage points.
+This pattern is consistent across models and prompting strategies, with the gap widening for smaller models.
+For instance, GPT-3.5 shows an even more dramatic difference: $86.28\%$ AIA vs $14.74\%$ MIA.
+This disparity suggests that while LLMs have become proficient at identifying correct reasoning paths, they struggle significantly when asked to recognize and articulate flawed reasoning.
+This challenge goes beyond simple counterfactual reasoning and points to a fundamental limitation in how these models process and evaluate logical structures.
+The ability to identify correct reasoning does not necessarily translate to the ability to pinpoint errors in incorrect reasoning, indicating a potential blind spot in the training or architecture of current LLMs.
+Impact of Model Size on MIA Performance
+The data reveals a clear correlation between model size and MIA performance.
+For the Math dataset with zero-shot CoT prompting, we see a progression from LLaMA3-8B ($27.05\%$ MIA) to LLaMA3-70B ($55.32\%$ MIA) to GPT-4o ($66.10\%$ MIA).
+This trend suggests that the ability to identify malgorithms is an emergent property that significantly improves with increased model capacity Wei et al.
+(2022a).
+However, it’s important to note that this improvement with model size is much more pronounced for MIA than for AIA.
+While GPT-4o outperforms LLaMA3-8B by $39.05$ percentage points in MIA, the difference in AIA is only $12.58$ percentage points ($95.65\%$ vs $83.07\%$).
+This disproportionate scaling indicates that malgorithm identification may require more complex reasoning capabilities that are only unlocked at larger model scales.
 Figure 1: An illustration of MIA performance across different Grades and DOK levels for Math and Reading question sets.
@@
 (best viewed in colors).
+Ineffectiveness of Few-Shot Learning
+The data clearly demonstrates that few-shot learning does not significantly improve MIA performance across models.
+For GPT-4o on the Math dataset, the MIA score with CoT prompting actually decreases slightly from $66.10\%$ (zero-shot) to $65.10\%$ (5-shot).
+LLaMA3-70B shows a similar trend, with MIA dropping from $55.32\%$ to $50.6\%$.
+Even in cases where there is a marginal improvement, the gains are negligible.
+For instance, LLaMA3-8B’s MIA increases from $27.05\%$ to $31.39\%$ with 5-shot CoT prompting - a difference of only $4.34$ percentage points.
+GPT-3.5 shows a similarly small improvement from $14.74\%$ to $17.06\%$.
+These results suggest that providing examples does not substantially enhance an LLM’s ability to identify malgorithms.
+This finding is particularly noteworthy given that few-shot learning often improves performance on other NLP tasks.
+The ineffectiveness of this approach for malgorithm identification underscores the unique challenges posed by this task and suggests that more sophisticated techniques may be needed to improve LLM performance in this area.
+4.3 Understanding CoT Prompting for MIA
+Our analysis of Chain-of-Thought (CoT) prompting reveals an unexpected ineffectiveness in improving Malgorithm Identification Accuracy (MIA), and in some cases, even leads to decreased performance compared to simple prompting.
+Inconsistent impact across subjects: For GPT-4o, CoT prompting shows negligible difference in Math MIA ($65.1\%$ vs $65.0\%$ for simple prompting), but significantly underperforms in Reading MIA ($68.6\%$ vs $82.8\%$ for simple prompting).
+Consistent underperformance of LLaMA-70B: LLaMA-70B demonstrates lower MIA prompting in both Reading ($58.4\%$ vs $60.0\%$) and Math ($48.3\%$ vs $50.6\%$) with CoT compared to simple prompting.
+Minimal impact on AIA: CoT prompting generally performs similarly to simple prompting for Algorithm Identification Accuracy (AIA) across models and subjects, indicating its ineffectiveness is specific to malgorithm identification.
+Figure 2: An illustration of MIA performance across five math content classes (Number & Operation, Algebra, Geometry & Measurement, Data Analysis, Probability) and two Reading content classes (Informational, Literature).
+These results challenge the assumption that explicit step-by-step reasoning always improves performance on complex tasks.
+The data suggests that CoT prompting, while effective for many NLP tasks, may actually hinder the model’s ability to identify flawed reasoning paths in the context of malgorithm identification.
+A plausible explanation for this counterintuitive finding lies in the nature of the training data typically used for LLMs.
+These models are predominantly exposed to examples of correct reasoning during pre-training and fine-tuning.
+Consequently, when CoT prompting guides the model through a structured thought process, it likely leverages these learned patterns of correct reasoning.
+This may inadvertently anchor the model to correct logical pathways, making it more challenging to deviate from these patterns and identify flawed reasoning.
+In essence, CoT prompting might be reinforcing the model’s bias towards correct reasoning, which is counterproductive when the task specifically requires identifying incorrect logical steps.
+5 Analysis: Educational Dimensions
 Performance across Grade Levels
@@
 Both models consistently show a decline in performance with increasing grade levels, with LLaMA3-70B exhibiting some unexpected variability, particularly excelling in Grade 8 of the reading dataset.
-These results highlight the increasing challenges in tackling malgorithm identification task as the questions becomes more complicated.
+These results highlight the increasing challenges in tackling malgorithm identification task as the questions becomes more complex.
 Performance across DOK
@@
 This observation suggests that larger models are better at processing structured, factual content, whereas smaller models are more efficient with narrative, context-rich material.
-Figure 2: An illustration of MIA performance for LLMs across five math content classifications (Number & Operation, Algebra, Geometry & Measurement, Data Analysis, Probability) and two reading content classifications (Informational Text, Literature).
-6 Related Works
-Counterfactual reasoning is a cognitive process that involves considering alternative scenarios to understand causality, which has gained increasing attentions with the advent of artificial intelligence.
-Pearl’s foundational work on causal inference [12] has established a theoretical framework for understanding counterfactuals, which has been applied to various domains, including fairness, interpretability, and robustness in machine learning [8, 17, 24].
-In the field of NLP, counterfactual reasoning is crucial for tasks that require understanding hypothetical scenarios and the causal relationships between events.
-However, most existing NLP models focus on predictive tasks rather than for hypothetical reasoning [3, 2, 6].
-While several datasets have been created to evaluate different aspects of reasoning in language models, relatively few focus explicitly on counterfactual reasoning.
-For instance, the bAbI dataset [28] includes tasks that require basic forms of logical reasoning, while the SQuAD dataset [15] focuses on reading comprehension.
-However, these datasets do not emphasize counterfactual scenarios.
-Recent efforts like the COPA dataset [16] and the CQA dataset [22] have introduced questions that require commonsense reasoning, which often involves understanding causal relationships.
-Nonetheless, these datasets are not explicitly designed to test counterfactual reasoning.
-To address this gap, we introduce MalAlgoQA, a novel dataset specifically designed to evaluate the counterfactual causal reasoning capabilities of LLMs through a pedagogical approach with multiple-choice question answering format.
-Applications of MalAlgoQA for personalized education: The use of LLMs in classrooms [20, 21], for instance, could profoundly influence the learning experiences of students.
+6 Implications for AI in Education
+The findings from our study on malgorithm identification have significant implications for the development and application of AI in educational contexts:
+Limitations in Understanding Student Misconceptions: The substantial performance gap between Algorithm Identification Accuracy (AIA) and Malgorithm Identification Accuracy (MIA) suggests that current LLMs are far more adept at recognizing correct reasoning than identifying flawed logic.
+For GPT-4o, we observed a drop from $95.65\%$ AIA to $66.10\%$ MIA on the Math dataset.
+This disparity indicates that AI tutoring systems based on these models may excel at confirming correct answers but struggle to diagnose and address student misconceptions effectively.
+This limitation could lead to missed opportunities for targeted intervention and personalized learning.
+Rethinking Training Approaches for Error Recognition The ineffectiveness of CoT prompting in improving MIA performance reveals a fundamental challenge in training LLMs to recognize and analyze errors.
+Rather than simply diversifying training data, this finding suggests a need for novel training paradigms that specifically target the skill of error identification Sonkar et al.
+(2024a).
+Future AIED systems may require specialized fine-tuning processes or architectural modifications that enable models to simultaneously reason about correct and incorrect logical pathways.
+This could involve developing new pre-training tasks focused on error detection, or creating model architectures that maintain separate representations for correct and flawed reasoning patterns.
+Such innovations could lead to AI tutors that are equally adept at reinforcing correct understanding and diagnosing misconceptions, significantly enhancing their effectiveness in personalized education.
+Rethinking Feedback Mechanisms: The challenges in malgorithm identification suggest that current AI models might provide incomplete or misleading feedback to students.
+AIED systems need to be designed with an awareness of these limitations, potentially incorporating human oversight or alternative assessment methods to ensure that student misconceptions are accurately identified and addressed.
+Implications for Automated Grading and Assessment: The performance discrepancy between AIA and MIA also raises concerns about the reliability of AI-driven automated grading systems, particularly for open-ended questions or problems requiring complex reasoning.
+Educational institutions considering the implementation of such systems should be aware of these limitations and consider alternate Sonkar et al.
+(2024b, d) or hybrid approaches with human review.
+Adaptive Difficulty Handling: The observed decline in MIA performance across increasing grade levels and DOK levels highlights the need for AIED systems that can adapt to varying levels of question complexity.
+Future AI tutors should be designed with mechanisms to maintain consistent performance across different difficulty levels, ensuring effective support for students as they progress through more challenging material.
+The difficulty LLMs face in identifying flawed reasoning points to a broader challenge in counterfactual reasoning.
+For AIED applications, this suggests a need to develop models with enhanced capabilities in exploring alternative scenarios and reasoning paths.
+Such improvements could lead to more sophisticated tutoring systems that can not only identify errors but also explain why certain approaches are incorrect and guide students towards correct understanding.
+By addressing these implications, AIED researchers can work towards not only improving the accuracy of student assessment but also enhancing the ability of AI systems to provide targeted and personalized support.
+7 Related Work
+Counterfactual reasoning, a critical cognitive process for understanding causality, has gained significant attention in machine learning (ML) and NLP research.
+Pearl’s seminal work on causal inference Pearl (2009) has been widely applied to address challenges in fairness Kusner et al.
+(2017), interpretability Wachter et al.
+(2018), and robustness Bareinboim and Pearl (2016) in ML.
+In NLP, counterfactual reasoning is important for tasks requiring comprehension of hypothetical scenarios.
+However, most existing NLP models primarily focus on predictive tasks rather than hypothetical reasoning Bommasani et al.
+(2021); Bender et al.
+(2021).
+Recent work has begun to address this gap developing specialized datasets and models for counterfactual reasoning Qin et al.
+(2019).
+Several datasets have been created to evaluate different aspects of reasoning, including counterfactual reasoning such as COPA dataset Roemmele et al.
+(2011), Counterfactual Story Rewriting task Qin et al.
+(2020), and CQA dataset Talmor et al.
+(2019).
+These efforts are complemented by development of new modeling approaches, such as causal LMs Feder et al.
+(2021), counterfactual data augmentation Kaushik et al.
+(2020), and generating counterfactual explanations Wu et al.
+(2021); Ross et al.
+(2021).
+Despite these advancements, challenges remain such as handling multi-step causal chains and generalizing across diverse domains Xu et al.
+(2023); Keith et al.
+(2020); Feder et al.
+(2022).
+8 Conclusion and Future Work
+This paper introduces MalAlgoQA, a novel dataset designed to evaluate the counterfactual reasoning capabilities of LLMs through a pedagogical lens.
+By proposing the Malgorithm Identification task and its associated metrics, we provide a new framework for assessing LLMs’ ability to recognize and analyze flawed reasoning patterns.
+Our findings reveal significant challenges in counterfactual reasoning for state-of-the-art LLMs.
+These results highlight the need for further research into enhancing LLMs’ counterfactual reasoning abilities, particularly for educational applications where understanding student misconceptions is crucial.
+Future work will focus on expanding the dataset, exploring novel training methodologies, and investigating architectural modifications to develop more cognitively-aligned language models capable of robust counterfactual reasoning.
+9 Limitations
+While our research sheds light on the counterfactual reasoning capabilities of LLMs, we acknowledge the inherent complexity and intricacy of the cognitive process may not be fully captured by our current evaluation framework.
+The MalAlgoQA dataset and the Malgorithm Identification task, although designed with careful consideration, are initial steps toward understanding the intricate aspect of cognition.
+The chain-of-thought (CoT) prompting technique is not a panacea and its effectiveness vary depending on the specific context and task.
+As research in this field progresses, we anticipate further refinements and enhancements to our methodology, contributing to a more comprehensive understanding of counterfactual reasoning in LLMs.
+10 Ethics and Risk
+The use of LLMs in classrooms, for instance, could profoundly influence the learning experiences of students.
 While LLMs can provide personalized learning support and instant feedback, their inability to accurately identify and address student misconceptions, as highlighted by our study, could potentially reinforce incorrect understanding or reasoning.
@@
 However, we acknowledge that this is a complex, multifaceted issue that requires continuous exploration, vigilance, and open dialogue among researchers, educators, policymakers, and society at large.
-7 Conclusion and Future Works
-This paper presents MalAlgoQA, a novel dataset designed to critically evaluate the counterfactual reasoning capabilities of LLMs through a pedagogical prespective.
-Additionally, we introduce the Malgorithm Identification task along with associated metrics for its evaluation.
-Our findings underscore the need for ongoing research and development the ability to identify malgorithms, particularly for educational applications where recognizing flawed reasoning processes is essential.
-Notably, our study also challenges the efficacy of the chain-of-thought prompting technique in enhancing counterfactual causal reasoning, advocating for further investigation into suitable prompting techniques for robust malgorithm identification.
-In the future, we plan to expand the MalAlgoQA dataset to cover a broader range of subjects and reasoning complexities.
-We also aim to explore different training methodologies and architectural modifications to LLMs that may enhance their ability to engage in counterfactual causal reasoning, ultimately contributing to the development of more cognitively-aligned language models.
 Acknowledgments
-This work was supported by NSF grants 1842378, ONR grant N0014-20-1-2534, AFOSR grant FA9550-22-1-0060, and a Vannevar Bush Faculty Fellowship, ONR grant N00014-18-1-2047.
-Appendix A Additional Analysis
-A.1 Grade vs DOK for different content classifications
-Figure 3: An image of MIA performance across different Grades and DOK levels for various Content Classifications for Math question sets.
-The top number represents MIA performance and the bottom number represents question counts.
-The data analysis and probability classifications are combined together for illustration.
-(best viewed in colors).
-Figure 4: An image of MIA performance across different Grades and DOK levels for two Content Classifications for Reading question sets.
-The top number represents MIA performance and the bottom number represents question counts (best viewed in colors).
-Figures 3 and 4 illustrate the relationships between grade levels, DOK for different content classifications in math and reading question sets.
-In mathematics, Number & Operation and Algebra questions demonstrate relatively consistent performance, with higher DOK levels often showing better results.
-For instance, MIA performance for Number & Operation peaks in lower grades and declines in higher grades, particularly Grade 7.
-This trend suggests increasing complexity in numerical operations as grade levels increases.
-On the other hand, Geometry & Measurements, Data Analysis and Probability exhibit high variability, with strong MIA performance in some grades but gaps in others.
-In reading, performance for Informational Text is generally consistent across grades, peaking in Grade 8, with lower DOK levels typically yielding higher MIA performance.
-Literature shows high variability, reflecting the varying complexity of literary analysis tasks.
-Notably, content classifications in mathematics show more consistent patterns across grades compared to reading ones.
-A.2 Performance across Grade level and Content Classifications
-Figure 5: An illustration of MIA performance across different Grades and Content Classifications for Math and Reading question sets.
-The top number represents MIA performance and the bottom number represents question counts (best viewed in colors).
-Figure 5 shows the relationship between grade levels and content classifications for math and reading question sets.
-In mathematics question set, Algebra demonstrates relatively consistent performance across all grades, with the number of algebra questions increasing significantly in higher grades, peaking at 191 data points in Grade 8.
-Number & Operation exhibits a gradual decline in MIA performance from lower to higher grades.
-In contrast, Data Analysis and Probability, Geometry & Measurement classifications show varied MIA performance across grade levels.
-For instance, Data Analysis and Probability demonstrates exceptionally high performance in Grade 6 (81.82), albeit with relatively few data points (22).
-In reading question set, MIA performance for both Informational Text and Literature classifications fluctuates across grades without much correlation.
-Informational Text achieves a significant peak in MIA performance for Grade 8, while Literature questions in Grade 5 demonstrate exceptionally high performance.
-This variability could be attributed to differences in text complexity or question difficulty at different grade levels for the Malgorithm Identification problem.
-A.3 Performance across Questions
-Figure 6: The distribution of GPT4-o performance in identifying malgorithms across question levels for math and reading question sets.
-The x-axis represents the number of correctly identified malgorithms per question, ranging from 0 to 3.
-We can find the performance of LLMs in identifying malgorithms across each question in both mathematics and reading comprehension sets in Figure 6.
-The number of questions increases as we progress from 0 to 3 correctly identified malgorithms per question.
-This pattern suggests that LLMs are more likely to identify multiple malgorithms correctly rather than missing all or most of them.
-However, despite this positive trend, it is crucial to note that LLMs’ overall performance on this task remains suboptimal.
-The presence of the significant number of questions at the 0 and 1 levels across both domains indicates room for improvement.
-In mathematics, the combined number of questions with 0 or 1 correctly identified malgorithms is comparable to those with 3 correct identifications.
-Similarly, in reading comprehension, while there is a clearer upward trend, the number of questions with 0 or 1 correct identifications remains considerable.
-These findings underscore the need for further research of LLMs’ counterfactual reasoning capabilities in malgorithm identification across various question types and subject areas.
-Appendix B Examples of Mathematics dataset
-We provide five examples, one per content classifications, of question, choices and their corresponding rationales in Table 8 and 9.
+This work was supported by NSF grant 1842378, ONR grant N0014-20-1-2534, AFOSR grant FA9550-22-1-0060, a Vannevar Bush Faculty Fellowship, and ONR grant N00014-18-1-2047.
+Appendix A Appendix
+Examples of MalAlgoQA dataset
+We provide five math examples, one per content classifications, of question, choices and their corresponding rationales in Table 8 and 9.
 Math Question Set Example 1:
@@
 Table 9: Examples of Mathematics question set.
-Appendix C Examples of Reading Comprehension dataset
-We provide two examples of passage, question, choices and their corresponding rationales in Table 10 and 11.
-Reading Question Set Example 1:
-Content Classification: Literature
-Passage: Blueberries
-1 You ought to have seen what I saw on my way
-2 To the village, through Pattersons pasture to-day:
-3 Blueberries as big as the end of your thumb,
-4 Real sky-blue, and heavy, and ready to drum
-5 In the cavernous pail of the first one to come!
-6 And all ripe together, not some of them green
-7 And some of them ripe! You ought to have seen!
-Question: In lines 1-7, the speaker expresses great surprise about the ...?
-Answer and Rationales:
-A: large pails of blueberries.
-Rationale A: Some readers might find this option plausible because the speaker mentions the cavernous pail of the first one to come" (line 5).
-However, this mention of the pail is referring to the pail into which the lucky picker who first stumbles across this crop will put the Blueberries as big as the end of your thumb" (line 3).
-The speaker does not express surprise that pickers are bringing large pails.
-B: number of blueberry pickers.
-Rationale B: Some readers might choose this option because the phrase the first one to come" in line 5 refers to a blueberry picker.
-However, although the speaker is predicting success for anyone who comes to pick the blueberries, the speaker is not expressing surprise that there are no pickers yet.
-Based on the speakers description of how big, blue, and ripe the berries are (as big as the end of your thumb, Real sky-blue, and heavy, and ready" (lines 3 and 4)), the berries are undisturbed, and no one has discovered them except for the speaker.
-C: ripeness of the blueberries.
-Rationale C: The speaker describes big as the end of your thumb" (line 3) blueberries that are Real sky-blue, and heavy," (line 4) and all ripe together" (line 6).
-The speaker repeats, You ought to have seen" in lines 1 and 7 to convey surprise.
-D: size of new blueberry bushes.
-Rationale D: This option might seem plausible because the size of the blueberries and their bushes are described.
-The first speaker draws attention to size when describing the blueberries as big as the end of your thumb" (line 3).
-In line 13, the second speaker exclaims, Why, there hasnt been time for the bushes to grow." This line implies that the bushes must be small, so it is surprising that the blueberries they produce are numerous and as big as the end of your thumb." Although lines 1-7 address the blueberries, the speaker does not mention the bushes on which they grow in those lines.
-Table 10: Example 1 of Reading question set.
-Reading Question Set Example 2:
-Content Classification: Informational Text
-Passage: Creating Codes: Grace Hopper and the Computer Compiler
-1 The computer screen freezes.
-No matter what or how often you click, the hourglass or rotating circle (that symbol that insists, Im working on it!") spins away.
-You want to shake the screen.
-Someone next to you says, I think you found a bug."
-Question: What is the main purpose of paragraph 1?
-Answer and Rationales:
-A: To demonstrate the importance of the subject
-Rationale A: Although some readers may conclude that the frozen computer screen in paragraph 1 demonstrates how important computer programming is currently, the paragraph describes a familiar feeling of frustration rather than detailing the more serious consequences that could result from programming mistakes.
-B: To help readers make a connection with the topic
-Rationale B: The paragraph describes a person facing a frozen computer screen that many readers have probably experienced in order to help readers make a personal connection to the topic.
-By describing this common situation, the readers interest is engaged.
-So, when the article continues on to discuss Hoppers contribution to programming and how she helped solve some of the early computer bugs, readers feel more connected with the topic.
-C: To give background information about the subject
-Rationale C: Although introductions often provide background information, and some readers may conclude that explaining a computer bug is background information for the article, the topic of the article focuses not on computer bugs but on Hoppers contributions to computer programming.
-D: To provide readers with a brief summary of the topic
-Rationale D: Although the main topic of the article is Hoppers contribution to computer programming and paragraph 1 mentions computers, paragraph 1 describes a common situation in order to raise interest among readers.
-It describes a person having computer problems rather than summarizing Hoppers contributions to computer programming.
-Table 11: Example 2 of Reading question set